what is the equation and the solution this word problem?
An advertisement for grape juice claims their juice contains 10% grape juice. How much pure grape juice would have to be added to 5 quarts of the drink to obtain a mixture of 40% grape juice.

Answer by RAY100(1637)

Let x = amt added (in quarts)
.
5 * 10% + x* 100% = (5+x) * 40%
.
5*.1 +x*1 =(5+x) *.4
.
.5 +x = 2 +.4x
.
.6x = 1.5
.
x= 2.5 quarts
.
check
.1*5 +1*2.5 = 7.5 *.4
.
.5 +2.5 = 3,,,,,,ok
.